# Is the panel fair? Evaluating panel compositions through network analysis. The case of research assessments in Italy

### v2

@@ 8 Discussion and conclusions
Research evaluation and especially administrative massive research evaluation such as the British REF have gained a central role in university and research policies, progressively reinforcing their function in gatekeeping, filtering, and legitimating knowledge. Research assessment is usually governed by panels of experts managing the evaluation processes. European Peer Review Guide of the European Science Foundation recommend that in the composition of panels “all affected parties” are fairly represented. The British REF and by the Italian research assessments formally adopted this fair representation. However, it is not easy to define who the affected parties are and what dimensions to take into account in defining their fair representation. The observable characteristics of scholars, such as research field, gender, affiliation and geography, seems not to be sufficient, in particular if a discipline is characterized by the coexistence of many schools of thought with different approaches, methodologies and policy recipes. This is precisely the issue pointed out for economics in the case of the research assessments in the United Kingdom and in Italy.
This paper proposes an empirical strategy for exploring if, despite the formal respect of the fair composition of panels in terms of easily observable characteristics of members such as gender and current university affiliation, it is possible to individuate hidden connections between members such that the panel composition could be considered as unfair. Three main type of connections are defined. The first is the direct or indirect collaboration of panellists in writing articles: a disproportionate diffusion of co-authorships among the members of a panel may indicate that its composition is restricted to people who have personal ties and possibly theoretical or methodological similarities. The second type of connections can be revealed by observing the set of scholarly journals where panellists published: a concentration of panel publication in a relatively restricted set of journals can be considered as an indicator of the lack of intellectual diversity in the panel. The third type of connections can be revealed by looking at whether restricted members have studied in a restricted set of institutions, whether they have had past or current affiliations in a restricted set of institutions, whether they have contributed to a restricted set of magazines, newspapers and blogs. Also in this case the existence of hidden connections can be considered as an indicator of limited intellectual and social diversity in the panel.
@@
The second objection was argued in an official document [ANVUR, 2012] in which ANVUR responded to early criticism of the 2004-2010 panel composition. The objection is that the panel “is a group of scholars with a high scientific profile and diversified in terms of skills and geographical origin, in full compliance with the criteria for the selection of panels published on the ANVUR website” [ANVUR, 2012]. This objection is valid if the analysis is limited to the actual, i.e. at the moment of panel formation, diversities in terms of geography and affiliations. This paper definitively documents that the diversification disappears when considering multiple affiliations, rather than just the the current one, and the academic careers of panel members.
@@
The third objection is also argued in the official document cited above. It takes seriously the question of the strong co-authorship links among the members of the panel 2004-2010, and it is based on the comparison of the co-authorship network generated by the panel with a control group built by considering the top-20 or top-50 Italian economists in Repec database (www.repec.org). The arguments sounds as follows: the coauthorship network linking the top-50 Italian economists and the president of the panel is similar to the one linking panel members. In Italy, in economics, all the scholars of “high scientific profile” collaborate. It is possible to counter-argue that the choice of the top-20 or top-50 italian economists in Repec is not really a control group, but an ad hoc choice. First of all: the call for panellists was open worldwide, hence it is not correct to limit the analysis to Italian economists. Moreover, Repec does not even represents a complete sample of Italian economists since it “is based on a limited sample of the research output in Economics and Finance. Only material catalogued in RePEc is considered. [..] Thus, this list is by no means based on a complete sample” [RePEc, 2017]. Finally, Repec ranking is based on bibliometrics and it is therefore biased towards gender, multidisciplinary methods, and any research orientations pursued by a minority of researchers in their respective disciplines [Corsi et al., 2019].
@@
Once documented that the composition of of the first two panels is unfair, it follows that also the results of the research assessments should be considered as unfair. In other words, procedural unfairness in panel composition determines the general unfairness of the research assessment, and unfair results. The notion of “unfair” results should be accurately distinguished from “biased” results. If we take seriously the central tenet that a fair panel composition determines not only a just procedure, but also substantively better decisions, some interesting questions arise: to what extent did procedural unfairness result in biased evaluations? How did unfair panel composition result in distortions in the research evaluation results? What were the mechanisms that translated the unfair composition into bias in the evaluations? What were the effect of excluding from the panels varieties of research experiences?
@@
More generally, our work shows the importance of a careful evaluation of panel composition in administrative research evaluation procedures. An unfair composition may be the result of a careless design: for example, a procedure involving a public call for panellists and then a random draw may not ensure a fair cross-sectional representation of the communities of scholars to be evaluated. But an unfair composition may also be the result of the capture of the regulator-evaluator by some segments of the scholarly community. At the other extreme, an unfair composition may be the result of the strategy of the regulator, who is able to predetermine the results of the evaluation by a suitable choice of panellists.
The results of our work show that it is possible to develop a research strategy that can verify the degree of fairness in panel composition by gathering information to represent their hidden connections. This search strategy can be appropriately adapted to different contexts for verifying the fairness of panel compositions in evaluation procedures.

### v3

@@ 8 Discussion and conclusions
Research evaluation and especially massive research evaluation such as the British REF have gained a central role in university and research policies, progressively reinforcing their function in gatekeeping, filtering, and legitimating knowledge. Moreover, research evaluation is a type of administrative process and as such it is subjected to specific rules and requirements, among which procedural fairness. Procedural fairness require that the panels of experts managing the evaluation processes have a fair composition. European Peer Review Guide of the European Science Foundation recommend that in the composition of panels “all affected parties” are fairly represented. The British REF and the Italian research assessments are formally bound to a fair composition of panels . However, it is not easy to define who the affected parties are and what dimensions to take into account in defining their fair representation. Some characteristics of scholars, such as research field, gender, affiliation and geography are easily verifiable. The analysis of these observable characteristics cannot be sufficient. Cognitive particularism or cronyism can arise if the panel is composed of scholars with a common view of their field, are linked by friendship or have a common view of scientific eminence [Cole et al., 1979]. This problem possibly arises in the evaluation of a discipline characterized by the coexistence of many schools of thought with different approaches, methodologies and policy recipes. This is precisely the problem that literature has highlighted for economics in the case of the research evaluation in the United Kingdom and in Italy.
This paper proposes an empirical strategy to test whether, despite the formal adherence to fair panel composition in terms of easily observable member characteristics such as gender and current university affiliation, it is possible to identify hidden connections between members such that panel composition could be considered unfair. Three main type of connections are defined. The first is the direct or indirect collaboration of panellists in writing articles: a disproportionate diffusion of co-authorships among the members of a panel may indicate that its composition is restricted to people who have personal ties and possibly theoretical or methodological similarities. The second type of connections can be revealed by observing the set of scholarly journals where panellists published: a concentration of panel publication in a relatively restricted set of journals can be considered as an indicator of the lack of intellectual diversity in the panel. The third type of connections can be revealed by looking at whether restricted members have studied in a restricted set of institutions, whether they have had past or current affiliations in a restricted set of institutions, whether they have contributed to a restricted set of magazines, newspapers and blogs. Also in this case the existence of hidden connections can be considered as an indicator of limited intellectual and social diversity in the panel.
@@
The second objection was argued in an official document [ANVUR, 2012] in which ANVUR responded to early criticism of the 2004-2010 panel composition. The objection is that the panel “is a group of scholars with a high scientific profile and diversified in terms of skills and geographical origin, in full compliance with the criteria for the selection of panels published on the ANVUR website” [ANVUR, 2012]. This objection is valid if the analysis is limited to the actual, i.e. at the moment of panel formation, diversities in terms of geography and affiliations. This paper definitively documents that the diversification disappears when considering multiple affiliations, rather than just the the current one, and the academic careers of panel members. Moreover, as noted by Cole et al. [1979, p. 34], this way of selecting panellists according to their scientific prominence may generate cognitive particularism: scholars who achieve “eminence tend to favor” others “who are similarly situated in the hierarchy of science”.
@@
The third objection is also argued in the official document cited above. It takes seriously the question of the strong co-authorship links among the members of the panel 2004-2010, and it is based on the comparison of the co-authorship network generated by the panel with a control group built by considering the top-20 or top-50 Italian economists in Repec database (www.repec.org). The arguments sounds as follows: the coauthorship network linking the top-50 Italian economists and the president of the panel is similar to the one linking panel members. In Italy, in economics, all the scholars of “high scientific profile” collaborate. It is possible to counter-argue by reiterating the above argument of the risk of cognitive particularism. Another specific counter-argument is that the choice of the top-20 or top-50 italian economists in Repec is not really a control group, but an ad hoc choice. First of all: the call for panellists was open worldwide, hence it is not correct to limit the analysis to Italian economists. Repec does not even represents a complete sample of Italian economists since it “is based on a limited sample of the research output in Economics and Finance. Only material catalogued in RePEc is considered. [..] Thus, this list is by no means based on a complete sample” [RePEc, 2017]. Finally, Repec ranking is based on bibliometrics and it is therefore biased towards gender, multidisciplinary methods, and any research orientations pursued by a minority of researchers in their respective disciplines [Corsi et al., 2019].
@@
Once documented that the composition of the first two panels is unfair, it follows that also the results of the research assessments should be considered as unfair. In other words, procedural unfairness in panel composition determines the general unfairness of the research assessment, and unfair results. The notion of “unfair” results should be accurately distinguished from “biased” results. If we take seriously the central tenet that a fair panel composition determines not only a just procedure, but also substantively better decisions, some interesting questions arise: to what extent did procedural unfairness result in biased evaluations? How did unfair panel composition result in distortions in the research evaluation results? What were the mechanisms that translated the unfair composition into bias in the evaluations? What were the effect of excluding from the panels varieties of research experiences?
@@
The main lesson to be learned from this work is that the fairness of panel composition is a key issue for the design of fair research evaluation procedures. The adoption of the prevailing recommendations regarding composition in terms of members’ gender, age, affiliation, or research field are important and easily verifiable. However, they are not sufficient to guarantee a fair composition of panels in terms of scholarly thinking, background, or policy orientation. Hence, these recommendations are not necessarily appropriate to prevent the emergence of cognitive particularism in research evaluation. The respect of formal requirements of the compositions in terms of gender, age, and so on does not necessarily prevent an unfair composition due to a careless design of the procedure of appointment: for example, a procedure involving a public call for panellists and then a random draw may not ensure a fair representation of “all affected parties”. Moreover, it does not prevent an unfair composition due to the capture of the regulator by some segments of the scholarly community. It does not even prevent an unfair composition due to a strategy of the regulator, who is able to predetermine the results of the evaluation by a suitable choice of panellists.
It is very difficult to draw indications about the rules that governments, universities, or other institutions should apply for building fair panels in research evaluation. Surely, procedural fairness requires the complete transparency and verifiability of the criteria used for appointing members to a panel. For instance, in case the institution would apply a fair representation of “all affected parties”, it should clearly define both the population representing “all affected parties” and the way in which panellists are chosen from it. In this case, randomization may be an eligible criterion.
The issue of designing fair panels has been predominantly addressed as a problem of fair representation by defining thresholds or quotas for observable characteristics of panelists, such as gender, age, research field. Building panel by respecting these thresholds or quotas poses feasibility problems to the institutions. Adding other criteria and quotas to be met for guaranteeing a fair intellectual composition of panels would make panel appointment practically unfeasible. To this end, it would probably be more effective, though not decisive, to clearly define incompatibilities between members. This may be very easy in relations to some feature emerged in this paper: for instance, neither two co-authors nor a supervisor and a supervisee in the past can be both members of a panel. It is instead much more difficult to define incompatibilities in relation to policy orientations, or methodological and epistemological views. The case study presented here shows that it is possible to test the fairness of panel composition by observing social and intellectual connections among panel members. This test, suitably adapted to the different contexts, can be used by authorities in charge of appointing the panels to verify the fairness of their compositions before evaluation procedures begin. This test, suitably adapted, can be used also by independent observers to challenge the procedural fairness of research evaluations.